I need to solve this problem withour using the quadratic equation.
Solve 3x ( x+3) = 2 (5x+1)

Answer by stanbon(75887)

solve this problem without using the quadratic equation.
Solve 3x ( x+3) = 2 (5x+1)
-------------------
3x^2+9x = 10x+2
3x^2 -x -2 = 0
Factor:
(3x+2)(x-1) = 0
x = -2/3 ; x = 1
====================
Cheers,
StanH.